OpenStudy (anonymous):

solve -11=18+s --- 5

OpenStudy (anonymous):

-11*5 = 18+s -55 -18 = s -73 = s

OpenStudy (anonymous):

is it \[-11=\frac{18+s}{5}\]?

OpenStudy (anonymous):

if so harkirat has it

jhonyy9 (jhonyy9):

s=-73